# The Impact of Renewable Energy Communities in the Italian Day-Ahead Electricity Market: A Scenario Analysis

### v1

@@ 4.2.2 Sensitivity analysis: mixed scenarios
Fourth, our results for the policy scenario indicate that RECs, by enabling higher self-consumption, could also reduce the need for costly investments in grid infrastructure. This finding aligns with the conclusions by Fuentes González et al. (2022), who modeled the effect of energy communities on transmission grid costs in Chile and found substantial system-wide savings. Backe et al. (2022) came to similar conclusions regarding the expansion of projected transmission capacity of the grid in six EU countries by 2060. Studies by Boccard and Goetz (2025) and Sarfarazi et. al. (2020) reveal that RECs could reduce pressure on the distribution grid too. Nonetheless, this effect is contingent on the size of a PV system and REC member composition in the former case and pricing strategies and community BESS operation in the latter case. While our study is not explicitly focused on grid cost optimization, the observed market impacts also suggest that RECs may alleviate pressure on the grid by decreasing energy demand from the external grid.
@@
Fifth, nuanced PV sizing will be essential to fully realize the potential of REC deployment not only to reduce stress on the grid but also wholesale market prices. Boccard and Goertz (2025) found that the different sizing of PV systems in RECs could induce opposite impacts on the distribution system. For example, RECs with homogeneous profiles of residential prosumers, which install PV systems that cover 25% - 50% of their load, reduce power exchanges with the grid, whereas those who install PV systems that cover 50%–100% of their load increase power exchanges with the grid due to the mismatch between energy injecting and self-consuming. The latter situation could create reverse flows and thus force additional costs for the DSOs that eventually will be shifted onto the non-members of RECs. In turn, RECs with higher number of residential prosumers with heterogeneous profiles and RECs with public prosumers, which install PV systems that cover 25% of their load, increase power exchanges with the grid, while those who install PV systems that cover 50% - 100% of their load reduce power exchanges with the grid. This drastically opposite impacts on power exchanges hint at importance of modeling not only diverse PV system sizes but also diverse categories of prosumers, as we do. In our simulations, we used real-world averaged PV sizes of REC prosumers and producers in Italy (Figure 8) differentiated by categories and electricity zones. Similarly to Boccard and Goertz (2025), our future research could adopt sensitivity tests based on differentiating PV sizes, which could suggest insights related to the REC PV sizing and changes to the wholesale equilibrium. This could enrich the discussions about policies for the REC deployment in a system-efficient welfare-enhancing manner.
Finally, our work contributes to the broader policy debate on energy independence and consumer resilience in a time of extreme market volatility. The 2022 energy crisis and ongoing geopolitical tensions have reaffirmed the importance of local, decentralized energy systems. By simulating multiple realistic REC uptake scenarios, we are able to outline the range of possible benefits, while also acknowledging their limits. Even in our most optimistic scenario (sc_\_mix2.2027), the net reduction in wholesale market quantities remains moderate, suggesting that RECs are a necessary but not sufficient condition for deep decarbonization, thus requiring other structural measures and investments.
6 Conclusions
This paper fills a relevant gap in the literature, by combining bottom-up engineering estimations, an extensive review of REC deployment in the Italian setting, and an economic synthetic (counterfactual) approach based on real micro-data from market outcomes. By relying on this data-driven approach, our study delivers a novel methodological framework to assess the impacts of RECs on the electricity market, with an application to the Italian DA market.
@@ 6 Conclusions
Our results show that REC deployment generates non-negligible effects in specific zones of the market (notably North and Central-South), while its overall system-wide impact remains quite moderate. Seasonal and hourly patterns confirm that self-consumption is a key driver of economic benefits from REC deployment at macro scale. Specifically, the estimated impacts on market volumes lie within a narrow range, from a minimum of -0.19% (January, Policy Scenario) to a maximum of +1.16% (April, BU Scenario). While both BU and HW scenarios yield consistently positive but limited net effects (up to +1.16% and +0.89%, respectively), the Policy scenario reveals mixed evidence: the impacts remain close to zero in most represented months, but turn slightly negative in January, suggesting that higher REC self-consumption can, under certain conditions, reduce the volumes available in the wholesale market.
@@
Indeed, by extending the analysis to seasonal and intra-day variations, our results confirm that the magnitude of REC impacts varies substantially across months and hours. Under the BU scenario, April emerges as the most responsive period, as market volumes with RECs surpass volumes without RECs by up to 3% during weekdays, while July exhibits a steadier impact close to 1% during peak hours. Conversely, cold months display more muted effects, with January showing negative deviations in the policy scenario due to the dominance of self-consumption patterns. Main factors behind this trend are the lower solar irradiation and the greater heating needs supplied by electricity in colder months. Lastly, when increasing the self-consumption rates for all categories of prosumers (from 45% to 55%) in line with Italian policy targets for RECs (sc_\_mix2.2027), the results yield an overall average reduction of about -0.3% on the equilibrium volumes of the Italian DA market, which is on 0.11% greater than the average reduction of -0.19% in the 45% self-consumption scenario. This finding indicates that the increase in self-consumption rates for RECs deepens the reduction in equilibrium quantities.
This latter observation suggests that by including BESS in energy communities, even greater diminishing of wholesale equilibrium quantities can be attained. The discharge of the community-owned BESS to cover the needs of REC members may entail the great systemic benefits during peak hours. An expensive and polluting thermal generation could be offset by a clean and social energy. Yet, as discussed by many authors (Soini et al., 2020; Schick et al., 2022; Chen et al., 2023b), a particular modality of a BESS system (e.g., a community-owned BESS or individual BESS for each prosumer within a REC, etc.) would eventually decide the scale of this beneficial systemic impact, although could even become adverse under certain circumstances (Sarfarazi et al., 2020).
@@
Standalone systems are included in our REC modeling corresponding to the actual deployment status in Italy. The reducing effect of RECs’ self-consumption on market quantities also suggests that both the volumes of energy injected into the grid from renewable standalone plants and volumes from prosumers’ plants within a REC could be offset by the self-consumption of REC prosumers at the distribution system where a REC is located. This finding contributes to the debate on the effects of the renewable-dominated system with high-electrified energy demand (incl. heating, cooling and transport) on wholesale market outcomes, supporting Böttger & Härtel (2022), Backe et al. (2022) and Riaz et al. (2019) previous results, in which the authors state that the broad diffusion of distributed energy resources can reduce and smoothen aggregate demand, consequently, decreasing wholesale market prices.
@@
This paper is not exempt from limitations. First, the calibration of scenarios relies on publicly available data on REC uptake and profiles, standard self-consumption rates, and the announced policy targets by 2027. Second, behavioral aspects and demand-side management responses are not fully captured, although they may play a decisive role in the future and in dynamic adjustments in demand profiles, mirroring peculiar weather conditions and real-time modifications. Third, the projected impact of REC deployment is evaluated under the assumption that the capacities of other RES technologies remain fixed. This allows us to isolate the marginal contribution of the REC deployment to market equilibrium, although the actual system impacts could differ if broader RES additions were accounted for. Fourth, we disregard consumers who participate in RECs, although theoretically they might amend its consumption profiles to maximize the incentive for ”shared energy” when joining RECs. However, as stated in Section 1, this degree of coordination has not yet been implemented in the Italian energy communities, moreover, at a substantial scale. Despite such limitations, our methodology has a wide replication potential. The synthetic counterfactual approach can be adapted to other EU (and non-EU) electricity markets to assess the role of RECs (or other distributed generation sources) on day-ahead market equilibrium. This would allow for a broader comparative insight into how local RECs can reshape aggregate electricity demand and supply, grid requirements, and market resilience.
@@
Finally, our findings provide timely inputs for energy and environmental policy-making. The diffusion of RECs could foster energy independence and price stability, particularly in current times of geopolitical uncertainty and high energy costs, as highlighted during the 2022 energy crisis. However, their legacy will depend on complementary measures: faster permitting procedures, harmonized governance frameworks, stronger support for storage and demand-side flexibility, and access to financing resources. Italian government extended a deadline that grants €2.2 billion in subsidies for new installations within RECs to six additional months, with a possible extension (Camera dei deputati, 2025). This optimistic signal suggests that actual REC deployment in Italy by December 2027 will not only match our ”Half-Way” scenario but also make the 5 GW ”Policy scenario” more attainable.
@@ 7.1 Appendix A
7.2 Appendix B

### v2

@@ 5 Discussion
Fourth, our results for the policy scenario indicate that RECs, by enabling higher self-consumption, could also reduce the need for costly investments in grid infrastructure. This finding aligns with the conclusions by Fuentes González et al. (2022), who modeled the effect of energy communities on transmission grid costs in Chile and found substantial system-wide savings. Backe et al. (2022) came to similar conclusions regarding the expansion of projected transmission capacity of the grid in six EU countries by 2060. Studies by Boccard and Goetz (2025) and Sarfarazi et. al. (2020) reveal that RECs could reduce pressure on the distribution grid too. Nonetheless, this positive effect is contingent upon the size of a PV system and composition of REC members in the former study and pricing strategies and community BESS operation in the latter study. While our study does not explicitly focus on optimization of grid costs, the observed market impacts also suggest that RECs may alleviate pressure on the grid.
@@
Fifth, nuanced PV sizing will be essential to fully realize the potential of REC deployment not only to reduce stress on the grid but also wholesale market prices. Boccard and Goertz (2025) found that the different sizing of PV systems in RECs could induce opposite impacts on the distribution system. For example, RECs with homogeneous profiles of residential prosumers, which install PV systems that cover 25% - 50% of their load, reduce power exchanges with the grid, whereas those who install PV systems that cover 50%–100% of their load increase power exchanges with the grid due to the mismatch between energy injecting and self-consuming. The latter situation could create reverse flows and thus force additional costs for the DSOs that eventually will be shifted onto the non-members of RECs. In contrast, RECs with greater number of residential heterogeneous prosumers and RECs with public prosumers, which install PV systems that cover 25% of their load, increase power exchanges with the grid, while those who install PV systems that cover 50% - 100% of their load reduce power exchanges with the grid. This drastically opposite impacts on power exchanges hint at importance of modeling not only diverse PV system sizes but also diverse categories of prosumers, as we do. In our simulations, we used real-world averaged PV sizes of REC prosumers and producers in Italy (Figure 10) differentiated by categories and electricity zones.
6 Conclusions and Policy Implications
@@
Our results show that REC deployment generates non-negligible effects in specific zones of the market (notably North and Central-South), while its overall system-wide impact remains quite moderate. Seasonal and hourly patterns confirm that self-consumption is a key driver of economic benefits from REC deployment at macro scale. Specifically, the estimated impacts on market volumes lie within a narrow range, from a minimum of -0.19% (January, Policy Scenario) to a maximum of +1.16% (April, BU Scenario). While both BU and HW scenarios yield consistently positive but limited net effects (up to +1.16% and +0.89%, respectively), the Policy scenario reveals mixed evidence: the impacts remain close to zero in most of the represented months, but turn slightly negative in January, suggesting that higher REC self-consumption can, under certain conditions, reduce the volumes available in the wholesale market. This observation becomes especially relevant if looking at the peak/off-peak dynamics. The reduction in equilibrium volumes during off-peak hours due to institutional prosumers within RECs (public, SMEs, NPOs) can reduce the need for RES curtailment while still having additional RES capacity deployed. On the other side, the increase in volumes during peak hours in spring and summer reduce the need for expensive thermal peaking plants. In contrast, when the increase in volumes occurs during off-peak hours, such as in scenarios with smaller capacity deployed (BU and HW), RECs adversely affect the market outcomes.
@@
Indeed, by extending the analysis to seasonal and intra-day variations, our results confirm that the magnitude of REC impacts varies substantially across months and hours. Under the BU scenario, April emerges as the most responsive period, as market volumes with RECs surpass volumes without RECs by up to 3% during weekdays, while July exhibits a steadier impact close to 1% during peak hours. Conversely, cold months display more muted effects. The main factors behind this trend are the lower solar irradiation and the greater heating needs supplied by electricity during cold months. Lastly, when increasing the self-consumption rates for all categories of prosumers (from 45% to 55%) in line with Italian policy targets for RECs (sc55.2027), the results yield an overall average reduction of about -0.3% on the equilibrium volumes of the Italian DA market, which is on 0.11% greater than the average reduction of -0.19% in the 45% self-consumption scenario. This finding indicates that the increase in self-consumption rates for RECs deepens the reduction in equilibrium quantities.
@@
Standalone systems are included in our REC modeling corresponding to the actual deployment status in Italy. The reducing effect of RECs’ self-consumption on zonal market quantities also suggests that both the volumes of energy injected into the grid from renewable standalone plants and volumes from prosumers’ plants within a REC could be offset by the self-consumption of REC prosumers. This finding contributes to the debate on the effects of the renewable-dominated system with high-electrified energy demand (incl. heating, cooling and transport) on wholesale market outcomes (Böttger and Härtel, 2022; Backe et al., 2022; Riaz et al., 2019).
Based on these insights, we outline several policy recommendations:
  1. 1.
    The greater the capacity deployed by RECs, the more pronounced the impact of institutional members (public entities, small and medium enterprises, and non-profit organizations), due to their traditionally higher volumes of self-consumed energy during off-peak hours compared with residential members (Figures 4, 5). The current Italian policy incentivizes smaller-capacity installations over larger ones by applying rigid capacity thresholds (see TIP in Figure 1 and Appendix 7.2), irrespectively of a prosumer category. However, rather than injecting energy into the grid during off-peak hours, rooftop installations that cover the loads of institutional members can greatly benefit the system. Therefore, the level of incentivization could be tailored to the extent of PV coverage of a building’s loads, instead of relying on rigid capacity thresholds. Such a measure could yield additional welfare gains as the electrification of public and industrial heating accelerates.
  2. 2.
    Institutional REC members inject more energy into the grid during peak hours in spring and summer than residential members, who tend to self-consume more energy during this period. Consequently, institutional REC members help meet peak demand in the system. In addition to public prosumers and non-profit organizations, small and medium-sized enterprises also exhibit load profiles that differ from those of residential prosumers. However, the Italian incentive scheme for shared energy does not make participation in RECs economically attractive for SMEs (see Koltunov et al. (2025)). The 50% reduction in the premium tariff for SMEs that benefit from capital subsidies constitutes one of the main obstacles. If such restrictive rules are relaxed, system-wide benefits from new RECs could increase.
  3. 3.
    According to the logic of Italian policy (MASE, 2023) and regulation (GSE S.p.A., 2025a; Schiavo et al., 2022), the economic systemic rationale behind RECs lies in the avoided use of the grid resulting from the simultaneous local injection and consumption of energy. We also demonstrate that reductions in market equilibrium volumes—particularly pronounced during the cold months—could help the system avoid RES curtailment when greater REC-driven capacity is deployed (i.e., 5 GW or more). This additional systemic rationale should be recognized by policy and thus valued/remunerated accordingly.
  4. 4.
    The injection of energy from both standalone facilities and rooftop solar installations can be offset by local self-consumption at the zonal scale. Since feed-in tariffs and feed-in premiums for large-scale renewables have been progressively phased out in the EU, the incentive for energy communities could also be used to support the rollout of standalone plants with capacities of up to 1 MW but in a more location-efficient and socially inclusive manner. Policymakers may therefore consider raising the maximum capacity threshold (i.e 1 MW) for individual REC configurations in order to enhance their economic viability for REC members as well as for third-party entities [^36] Participation in RECs is not open to large companies or companies whose primary activity belongs to energy sector. Therefore, utilities and ESCOs tend to participate in ”energy sharing” scheme as third-party entities. This means they cannot directly benefit from capital subsidy and from incentive on shared energy but can request lease payments if they own a renewable installation and/or charge configuration management fee to RECs.
  5. 5.
    By adopting battery storage within energy communities, even greater reductions in wholesale market equilibrium quantities can be achieved. The discharge of community-owned batteries to meet the needs of REC members generates substantial systemic benefits during peak hours. In this way, costly and polluting thermal generation can be displaced by clean and social energy. Still, as was discussed by many authors (Soini et al., 2020; Schick et al., 2022; Chen et al., 2023b; Sarfarazi et al., 2020), the specific modality of a storage system ultimately determines the scale and direction of this systemic impact. In Italy, capital subsidies are limited to 40% of the costs incurred by RECs, a design feature that discourages the installation of costly storage. To avoid additional public spending while still promoting systemic efficiency, policy could specifically incentivize integrated photovoltaic–battery systems in regions with excess renewable generation (e.g., Southern Italy).
  6. 6.
    RECs have historically been regarded as collective action initiatives created primarily by citizens and for citizens (e.g., energy cooperatives). Recently, Italian policy has also allowed institutional actors to participate in RECs. The load profiles of institutional prosumers enable greater systemic benefits from REC deployment—particularly when appropriately aggregated with residential loads—than those associated with multiple residential prosumers. Consequently, other countries may also consider permitting institutional actors to participate in RECs.
@@
This paper is not exempt from limitations. First, the calibration of scenarios relies on publicly available data on REC uptake and profiles, standard self-consumption rates, and the announced policy targets by 2027. Second, behavioral aspects and demand-side management responses are not fully captured, although they may play a decisive role in the future and in dynamic adjustments in demand profiles, mirroring peculiar weather conditions and real-time modifications. Third, the projected impact of REC deployment is evaluated under the assumption that the capacities of other RES technologies remain fixed. This allows us to isolate the marginal contribution of the REC deployment to market equilibrium, although the actual system impacts could differ if broader RES additions were accounted for. Fourth, we disregard consumers who participate in RECs, although theoretically they might amend its consumption profiles to maximize the incentive for ”shared energy” when joining RECs. However, as stated in Section 1, this degree of coordination has not yet been implemented in the Italian energy communities, moreover, at a substantial scale.
Despite limitations, our methodology has a wide replication potential. The synthetic counterfactual approach can be adapted to other EU (and non-EU) electricity markets to assess the role of RECs (or other heterogeneous distributed energy resources) on day-ahead market equilibrium. This would allow for a broader comparative insight into how local RECs can reshape the aggregate electricity demand and supply, price signals, grid requirements and market resilience. A robustness check under an “accelerated battery adoption scenario” could further refine our estimates, offering a valuable sensitivity benchmark for future policy analysis. Similarly to Boccard and Goertz (2025), our future research could also adopt sensitivity tests based on differentiating PV sizes, which could suggest insights related to PV sizing for energy communities and its consequences for changes in a wholesale equilibrium.
@@
Finally, our findings provide the timely inputs for energy and environmental policy-making. The diffusion of RECs could foster systemic efficiency, energy justice and independence, particularly in current times of geopolitical uncertainty and high energy costs, as highlighted during the 2022 energy crisis. However, their legacy will depend on complementary measures: efficient policy and regulation, harmonized governance frameworks, stronger support for storage and demand-side flexibility, and access to financing resources. The Italian government extended a deadline that grants €2.2 billion in subsidies for new installations within RECs to six additional months, with a possible extension (Camera dei deputati, 2025). This optimistic signal suggests that actual REC deployment in Italy by December 2027 will not only match our ”Half-Way” scenario but also make the 5 GW ”Policy scenario” more attainable.
@@ 7 Appendices
JARSC Jointly acting renewable self-consumers
7.2 Appendix B. Shared energy and cashback calculation
In Italian regulation, shared energy refers to the minimum amount between the total energy injected into the public grid by all EC members within a given hour and the total energy withdrawn (consumed) by all EC members within the same hour.
The generalized formula for calculating shared energy is:
$E_{\mathrm{sh}}(h)=\min\left\{\sum_{i=1}^{n}E^{\mathrm{in}}_{i,h}\;;\;\sum_{i=1}^{n}E^{\mathrm{out}}_{i,h}\right\}$ (24)
where:
$E_{\mathrm{sh}}$:
the shared energy,
$E^{\mathrm{out}}_{i,h}$:
the energy withdrawn from the grid in a given hour $h$ by the $i$-th member,
$E^{\mathrm{in}}_{i,h}$:
the energy fed into the grid in a given hour $h$ by the $i$-th member.
The cashback logics is given by ARERA (2022) and Technical Rules from GSE (2025a). It can be described as:
$\mathrm{Cashback}=E_{\mathrm{inj}}\,P_{\mathrm{sell}}\;+\;E_{\mathrm{sh}}\,TIP\;+\;V_{\mathrm{agus}}\;-\;\mathrm{GSE\;costs}$ (25)
where:
$P_{\mathrm{sell}}$:
price for which GSE is purchasing renewable energy from eligible subjects. Regulated by ARERA 280/07 and equivalent to the Prezzo Minimo Garantito (PMG).
TIP :
premium tariff for shared energy.
$V_{\mathrm{agus}}$:
valorization of avoided grid usage due to self-consumption.
GSE costs:
administrative costs incurred by GSE related to cashback calculation and remuneration.
It is critical to consider how distinct components of the cashback are calculated. The premium tariff itself is not a single value; the basic premium tariff for an EC consists of fixed and variable components:
[…]
where
  • 80-70-60 are fixed components of a premium tariff that according to CER Decree stands at 60–80 €/MWh depending on the asset capacity (if $P<200$ equals 80; if $200<P<600$ equals 70; if $P>600$ equals 60). PV systems in central and northern Italy get an extra premium of 4 and 10 €/MWh respectively.
  • $\max(0;\,180-P_{z})$ is a variable component with $P_{z}$ being the hourly zonal price of electricity. In the event of zonal prices being higher than 180 €/MWh the variable component is counted as 0 €/MWh. This is made to prevent excessive premium during the periods of wholesale price spikes.
  • 120-110-100 are components that vary by installed capacity. If $P>600$ equals 100 €/MWh, if $200<P<600$ equals 110 €/MWh, if $P>600$ equals 120 €/MWh. In the event of extremely low wholesale price $P_{z}$ this parameter prohibits exceeding the premium tariff beyond 100–120 €/MWh.
The valorization of avoided grid usage due to self-consumption Vagus includes two components: the reduction in grid costs due to self-consumption and the reduction in costs related to decreased grid losses from self-consumption. JARSCs are compensated for both components, whereas RECs receive compensation only for the reduction in grid costs due to self-consumption. This distinction arises because RECs still rely on the distribution grid for energy sharing, thereby incurring grid losses. In contrast, JARSCs operate within a single condominium, bypassing the distribution grid. The Vagus for JARSCs and RECs can be calculated using the following formulas:
$\text{JARSC: }V_{agus}=T+L$ (27)
$\text{REC: }V_{agus}=T$ (28)
where
  • $T$ is reduction of grid costs due to self-consumption
  • $L$ is reduction of costs related to decreased grid losses due to self-consumption
Reduction of grid costs due to self-consumption (T) is calculated (Schiavo et al., 2022) by formula:
$T=E_{sh}\times\bigl(TRASe+\max(BTAUm)\bigr)$ (29)
where
  • TRASe is a sum of transmission tariff for LV household consumers,
  • BTAUm is a distribution tariff for LV commercial and industrial consumers
Reduction of costs related to decreased grid losses due to self-consumption ($L$) is calculated (Schiavo et al., 2022) by formula:
$L=\sum_{i,h}\bigl(E_{sh}\times c_{agl}\times P_{z}\bigr)$ (30)
where
  • $c_{agl}$ is a coefficient for avoided grid losses of the voltage level to which generator/s are connected (2.6% for generators in LV grid, 1.2% for generators in MV grid),
  • $P_{z}$ is the hourly zonal energy price of the day-ahead market,
  • $E_{sh}$ is the energy shared by the $i$-th member
By identifying all components of the cashback, we can calculate the actual economic benefit for EC members:
$\textit{Benefit}=E_{sc}(P_{buy})+[E_{inj}(P_{sell})+E_{sh}(P_{r})+V_{agus}-GSE_{costs}]$ (31)
where
  • $E_{sc}$ is an amount of self-consumed energy by common loads for JARSCs and by prosumers for EC,
  • $P_{buy}$ is a retail price that a member of EC pays to its retailer.
7.3 Appendix C